# On an Optimal Stopping Problem with a Discontinuous Reward

### v3

@@ 3.2. Surrender Right - Non-Trivial Case
A right-continuous adapted process $X=\{X_{t}\}_{t\geq 0}$ defined on a filtered probability space $\{\Omega,\mathcal{F},\{\mathcal{F}_{t}\}_{t\geq 0},\mathbb{P}\}$ is said to be of class (D) if the family $\{X_{\tau},\tau\in\mathcal{T}_{0,\infty}\}$ is uniformly integrable.
@@
In light of Theorem 3.3, proving the existence of an optimal stopping time is simply a matter of checking that the discounted reward process is positive, optional and of class (D), and has upper semi-continuous trajectories.
@@
The discounted reward process $Z=\{Z_{t}\}_{0\leq t\leq T}$, with $Z_{t}=e^{-rt}\varphi(\tau,F_{t})$, is positive and adapted with upper semi-continuous trajectories. Moreover, it is dominated by an integrable non-negative random variable, $\sup_{0\leq t\leq T}e^{-rt}\varphi(t,F_{t})$, so it is of class (D). The result then follows from Theorem 3.3. ∎
@@
The modified reward function $g(t,x)x\vee h(t,\,x)$ is continuous since it is the maximum of two continuous functions. Indeed, $g(t,x)x$ is continuous by definition while the continuity of $h(t,\,x)$ follows from Theorem 3 of Veretennikov (1981), which is the analog of the Feynman-Kac Theorem (see, for instance, Karatzas and Shreve (1991), Theorem 5.7.6).
@@
the smallest optimal stopping time for (13), as per Theorem 3.3.
@@
$\qquad+\mathbb{E}\left[\mathbb{E}\left[e^{-r(T-t)}\max(G,F_{T}^{t,x})\mathbf{1}_{\{h(\widetilde{\tau},F_{\widetilde{\tau}}^{t,x})>g(\widetilde{\tau},F_{\widetilde{\tau}}^{t,x})F_{\widetilde{\tau}}^{t,x}\}}\Big{|}\mathcal{F}_{\widetilde{\tau}}\right]\right]$
@@
It is possible to show that $\bar{\tau}_{t}^{x}=\tau_{t}^{x}$ a.s., where $\tau_{t}^{x}$ is the smallest optimal stopping time for the original problem defined in Corollary 3.1. A comparison of the three stopping time is given in lemma 3.3 below.
@@
(1) By the definition of $\tau_{t}^{x}$ and the reward function $\varphi$, we have ${v(\tau_{t}^{x}(\omega),F^{t,x}_{\tau_{t}^{x}(\omega)}(\omega))=g(\tau_{t}^{x}(\omega),F^{t,x}_{\tau_{t}^{x}(\omega)}(\omega))F^{t,x}_{\tau_{t}^{x}(\omega)}(\omega)}$, and if $\widetilde{\tau}_{t}^{x}(\omega)=\tau_{t}^{x}(\omega)$, it follows that ${\widetilde{\tau}_{t}^{x}(\omega)=\tau_{t}^{x}(\omega)=\bar{\tau}_{t}^{x}(\omega).}$ From the definition of $\tau_{t}^{x}(\omega)$, we know that ${v(s,F^{t,x}_{s}(\omega))>g(s,F^{t,x}_{s}(\omega))F^{t,x}_{s}(\omega)}$ for all ${t\leq s\leq\tau^{x}_{t}(\omega).}$ Thus, for $\widetilde{\tau}_{t}^{x}(\omega)$ to be equal to $\tau_{t}^{x}(\omega)$, we need ${v(s,F^{t,x}_{s}(\omega))>h(s,F^{t,x}_{s}(\omega))}$, for all ${t\leq s\leq\tau^{x}_{t}(\omega)}.$ Now, suppose there exists some $s$ in the interval $[t,\tau_{t}^{x}(\omega)]$ such that ${v(s,F^{t,x}_{s}(\omega))=h(s,F^{t,x}_{s}(\omega))}.$ In this case, the two stopping times $\tau_{t}^{x}(\omega)$ and $\bar{\tau}_{t}^{x}(\omega)$ would differ. However, the equality ${v(s,F^{t,x}_{s}(\omega))=h(s,F^{t,x}_{s}(\omega))}$ would imply that the probability that the contract is exercised before maturity is nil, which contradicts our initial assumption that $\tau_{t}^{x}(\omega)<T$. Therefore, it must hold that ${v(s,F^{t,x}_{s}(\omega))>h(s,F^{t,x}_{s}(\omega))}$ for all ${t\leq s\leq\tau^{x}_{t}(\omega)}$, which leads to $\widetilde{\tau}_{t}^{x}(\omega)=\tau_{t}^{x}(\omega)=\bar{\tau}_{t}^{x}(\omega).$
@@
The intuition behind the equality between $\tau^{x}_{t}$ and $\bar{\tau}^{x}_{t}$ is that $\bar{\tau}^{x}_{t}$ depends on why the contract is surrendered in the problem with the continuous reward. Indeed, the strategy underlying $\widetilde{\tau}^{x}_{t}$ is to stop if the contract value is equal to either the surrender benefit or the expected present value of the maturity benefit. If the reason to stop is equality with the surrender benefit, then this optimal stopping time coincides with the strategy underlying the problem with the discontinuous reward function and all three stopping times are equal. However, if the reason to stop is that the value of the contract is equal to the EPV of the maturity benefit, then $\bar{\tau}^{x}_{t}$ is set to maturity $T$. In that case, it is also optimal to wait until maturity in the original problem, since the contract being worth the EPV of the maturity benefit means that it is never be optimal to surrender before maturity.
@@
In the next section, we present a condition on the fee and surrender charge functions under which $\widetilde{\tau}_{t}^{x}=\tau_{t}^{x}$ a.s. This result is particularly interesting since it provides a condition under which the two optimal stopping problems defined in (7) and (13) are equivalent. That is, the two problems have the same value function (Theorem 3.4), surrender region (Proposition 4.8), and optimal stopping time (Corollary 4.3).
@@ 4. Analytical study of the value function
The representation of the value function in terms of a continuous reward function simplifies the analysis of its regularity, which is presented in Section 4.1. However, additional work is required to assess its smoothness, which we discuss in Proposition 4.1 below. In Section 4.2, we establish the relationship between $v$, a free-boundary value problem, and a variational inequality. To do so, we use the continuity of the value function and follow the reasoning of Jacka (1991) and Lamberton (1998) in the context of American put option pricing. This allows us to derive, in Section 4.3, two other representations for the value function: the surrender premium representation which is analogous to the exercise premium representation in the American option pricing terminology, and the continuation premium representation. This representation is new to the literature on VA and American option pricing and allows us to characterize the (non-)emptiness of the surrender region in Section 4.4. Section 4.5 presents a condition on the fee and the surrender charge functions under which the optimal stopping problem with discontinuous reward function defined in (7) is equivalent to the one with the continuous reward function, in (13).
Assumptions 4.1 and 4.2 presented next concern the fee and the surrender function and are needed for several of the results of this section to hold. In Assumption 4.1 below, ${\mu:[0,T]\times\mathbb{R}_{+}\rightarrow\mathbb{R}}$ denotes the drift term of the account value process (4), that is, $\mu(t,x)=(r-C(t,x))x$.
@@
    For all $x$, the fee function $C(\cdot,x)$ is locally Hölder continuous.
@@ 4.1. Elementary Properties of the Value Function
The value function $v$ is continuous on $[0,T]\times\mathbb{R}_{+}$.
@@
For every $(t,\,x)\in E:=[0,\,T]\times\mathbb{R}_{+}$, the value function $v$ satisfies the following properties:
@@
The properties given in Lemma 4.1 follow from the definition of the contract and hold in a more general setting. They do not require Assumptions 4.1 and 4.2 to be satisfied.
@@
Proposition 4.1 below concerns the smoothness of the value function and is inspired by Lamberton (1998), Proposition 10.3.5 in the context of the American put under Black-Scholes. Our proof, which can be found in the Appendix, presents non-trivial differences compared to the American put case, which are due to the time-dependence of the value function.
@@ 4.2. Free Boundary Value Problem and Variational Inequality
While there exist significant differences between our problem and the well-studied pricing of an American option, the continuous reward representation of Theorem 3.4 and the previously established continuity of the value function allows us to apply classical results from optimal stopping theory. In this section, we use this strategy to establish the relationship between the value function $v$ associated with the original optimal stopping problem 19, a free-boundary value problem, and a variational inequality. Thus, the proofs of the results presented in this section are kept short; the reader is referred to the literature for more details.
The value function is supported on $[0,T]\times\mathbb{R}_{+}$, which is separated into the surrender region $\mathcal{S}$ and the continuation region $\mathcal{C}$, defined by
@@
respectively. It follows that $\mathcal{C}\cup\mathcal{S}=[0,T)\times\mathbb{R}_{+}$, since $v(t,x)\geq\varphi(t,x)$ for all $(t,x)\in[0,T]\times\mathbb{R}_{+}$. Furthermore, since $v$ and $\varphi$ are continuous on $[0,T)\times\mathbb{R}_{+}$ (see Theorem 4.1) , we have that the continuation region is open and the surrender region is closed. The exercise (or surrender) boundary is the boundary $\partial\mathcal{C}$ of $\mathcal{C}$.
To express the value function as the solution of a free-boundary problem, we define the second-order differential operator $\mathcal{L}_{t}$, for $0\leq t\leq T$, by
@@
Note that $L$ is the term on the right-hand side of (9). We give this general definition of $L$ since many of our results apply in a more general setting. However, under Assumption 4.2, which is assumed to hold throughout Section 4, $L$ becomes
$L(t)=g_{t}(t)-c(t)g(t).$
The next results confirm that $v(t,x)$ solves a free-boundary problem.
@@
The value function $v$ is the unique solution to the boundary value problem
@@
That is, any solution $v$ to (19) can be expressed by (7). In particular, the function $v\in C^{1,2}$ on $\mathcal{C}$.
@@
Once the continuity of the value function has been established (Theorem 4.1), the optimal stopping time has been shown to exist (Corollary 3.1), and the martingale property of the Snell envelope of the discounted reward process has been proven (see Theorem 3.3 (ii), the connection between partial differential equations and optimal stopping problems can be established assuming enough regularity of the coefficients in (4). A sufficient condition for the results of Friedman (1964) to hold is that the drift term in (4) is locally Hölder continuous, which is why Assumption 4.1 is needed in Theorem 4.2. The proof makes use of standard partial differential equation (PDE) results in solving the Dirichlet (or the first initial-boundary value) problem, see for instance Friedman (1964), Theorem 3.4.9.
The main difficulty in establishing the free-boundary value problem resides in proving the continuity of the value function (Theorem 4.1). Once this is done, the free-boundary value problem follows from standard arguments. Specifically, the continuity of the value function implies that the continuation region is open. Then, the Dirichlet problem (or the first initial-boundary value problem) can be posed in an open subset (typically a ball or a rectangle) of the continuation region. Since the boundary of the defined open subsets is sufficiently regular, standard PDE results guarantee that the Dirichlet problem admits a unique solution. Probabilistic arguments are then used to identify this solution as the value function.
For details of the proof in the context of the American put in the Black-Scholes setting, see Jacka (1991), Proposition 2.6 or Karatzas and Shreve (1998), Theorem 2.7.7, among others. For more general results, see Peskir and Shiryaev (2006), Sections 3.7.1 and 4.8.2 or Jacka and Lynn (1992), Proposition 3.1.
Assumption 4.1 does not give much information on the shape of the surrender region. However, we show in Section 4.4 (Theorem 4.3) that under further conditions on the fee and the surrender charge functions, the surrender region can be expressed in terms of the optimal surrender boundary $b:[0,T]\mapsto\mathbb{R}_{+}$. Then, the free-boundary value problem can be stated more explicitly.
@@
If $L(t)<0$ for all $t\in[0,T)$, the value function $v$ is the unique solution to the boundary value problem
@@
with $b(t):=\inf(\mathcal{S}_{t})$. That is, any solution $v$ to (19) can be expressed by (7). In particular, the function $v$ is $C^{1,2}$ on $\mathcal{C}$.
@@
For each $t\in[0,T)$, $L(t)<0$ is equivalent to $\frac{\,\mathrm{d}\ln g(t)}{\,\mathrm{d}t}<c(t)$.
@@
When $L(t)<0$ for all $t\in[0,T]$, we have from Corollary 4.1 that $v\in C^{1,2}$ on the continuation region, that is, when $(t,x)\in[0,T]\times(0,b(t))$. On the other hand, when $(t,x)\in[0,T]\times[b(t),\infty)$, we know that $v(t,x)=g(t)x$, so $v\in C^{1,2}$ on $\text{int}(\mathcal{S})$. What remains to assess is the smoothness of the value function across the boundary (i.e. when $x=b(t)$). When $L(t)\geq 0$ for all $t\in[0,T]$, an optimal stopping time for (7) is $T$, as per Proposition 3.1. Hence, $v(t,x)=h(t,x)$, the present value of the maturity benefit (5). The regularity of the value function then follows from the Feynman-Kac representation (see Karatzas and Shreve (1991), Theorem 5.7.6), which states that, under some regularity assumptions on the coefficients of (4), $h$ is the unique solution to a Cauchy problem, and such a solution is in $C^{1,2}([0,T)\times\mathbb{R}_{+})$.
Next, we discuss the link between the value function and a variational inequality. In Proposition 4.2 below, we establish a result that is a direct consequence of the process $J=\{J_{t}\}_{0\leq t\leq T}$, with $J_{t}:=e^{-rt}v(t,\,F_{t})$, being the Snell envelope of the discounted reward process $Z$, and thus a supermartingale by Theorem 3.3 (i).
@@
The value function $v$ defined in (7) satisfies
(21) $\mathcal{L}_{t}v+v_{t}-rv\leq 0$
for all $(t,x)\in(0,T)\times\mathbb{R}_{+}$, with all derivatives in the sense of distribution.
Proposition 4.2 is analogous to Proposition 10.3.7 of Lamberton (1998) in the context of the American put. For an introduction to weak derivatives and distribution theory see Evans (2010), Chapter 5, and Rudin (1991), Chapter 6.
The idea of the proof of Proposition 4.2 rests on the assumption that the value function is regular enough to apply Itô’s lemma to $J$ (which is usually not the case since $v$ may not be smooth enough across the boundary, see Remark 4.9). $\mathcal{L}_{t}v+v_{t}-rv$, being the drift term of $J$, must be non-positive since $J$ is a supermartingale. In the context of American put option pricing, a rigorous proof relying on distribution theory is given in Lamberton (1998), Proposition 10.3.7. Their arguments can easily be adapted to our problem, as long as the fee function does not depend on the account value.
@@
Proposition 4.2 holds in a more general setting, as it only requires Assumption 4.2 to be satisfied. In fact, its proof only needs the fee function to be independent of the account value; the condition on the surrender charge function in Assumption 4.2 can be relaxed.
From Theorem 4.2 and Proposition 4.2, we obtain that the value function $v$ solves a variational inequality.
Proposition 4.3.
Let Assumptions 4.1 and 4.2 hold. The value function $v$ defined in (7) is a solution to the variational inequality
(22) $\max\left\{\mathcal{L}_{t}v+v_{t}-rv,\varphi-v\right\}=0,$
with terminal condition $v(T,x)=\max(G,x)$ and all derivatives in the sense of distribution.
Remark 4.11.
Proposition 4.3 holds as long as Assumption 4.1 is satisfied and the fee function does not depend on $x$.
@@
The partial derivatives (in the sense of distribution) $\partial v/\partial x$, $\partial v/\partial t$ and $\partial^{2}v/\partial x^{2}$ are locally bounded, and $\partial v/\partial x$ is continuous on $[0,T)\times\mathbb{R}_{+}$.
@@
From Proposition 4.1, we have that the first order derivatives of the value function (in the sense of distribution) in $t$ and in $x$ are locally bounded. Using the convexity of $x\mapsto v(t,x)$ (Lemma 4.2), the local boundedness of the first order derivatives and (21), we can show that the second order derivative in $x$ is also locally bounded (see for instance Lamberton (1998), Theorem 10.3.8 or Jaillet et al. (1990), Theorem 3.6 for details). This last result is also known as the smooth fit condition in the American option terminology.
The continuity of $\frac{\partial v}{\partial x}$ on $[0,T)\times\mathbb{R}_{+}$ is obtained using the first part of the statement; any function $u(t,x)$ defined on $\mathbb{R}\times\mathbb{R}$ has a continuous first partial derivative in $x$ if $\partial u/\partial x$, $\partial u/\partial t$ and $\partial^{2}u/\partial x^{2}$ are bounded, see Ladyz̆enskaja et al. (1968), Chapter 2, Lemma 3.1. Indeed, Lipschitz-continuity in $x$ comes from the boundedness of $\partial^{2}U/\partial x^{2}$, and the boundedness of both $\partial u/\partial t$ and $\partial^{2}u/\partial x^{2}$ are used to show continuity in $t$ (see the proof of Corollary 10.3.10 of Lamberton (1998) for a sketch of the proof).

### v4

@@ 3.2. General Case
A right-continuous adapted process $X=\{X_{t}\}_{t\geq 0}$ defined on a filtered probability space $\{\Omega,\mathcal{F},\{\mathcal{F}_{t}\}_{t\geq 0},\mathbb{P}\}$ is said to be
  • of class (D) if $\{X_{\tau},\tau\in\mathcal{T}_{0,\infty}\}$ is uniformly integrable;
  • of class (DL) if for each $t\geq 0$, $\{X_{\tau},\tau\in\mathcal{T}_{0,t}\}$ is uniformly integrable.
@@
In light of Theorem 3.3, since our discounted reward process is positive and adapted with upper semi-continuous trajectories, showing the existence of an optimal stopping time is simply a matter of verifying that $Z$ is of class (D).
@@
Since $Z=\{Z_{t}\}_{0\leq t\leq T}$, with $Z_{t}=e^{-rt}\varphi(t,F_{t})$, is dominated by an integrable non-negative random variable, $\sup_{0\leq t\leq T}e^{-rt}\varphi(t,F_{t})$, it is of class (D). The result then follows from Theorem 3.3. ∎
@@
The modified reward function $g(t,x)x\vee h(t,\,x)$ is continuous since it is the maximum of two continuous functions. Indeed, $g(t,x)x$ is continuous by definition while the continuity of h(t,x)h(t,\,x) follows from Theorem 3 of Veretennikov (1981), which is the analog of the Feynman-Kac Theorem (see, for instance, Karatzas and Shreve (1991), Theorem 5.7.6).
@@
the smallest optimal stopping time for (13), as per Theorem 3.3(iii).
@@
[…]
@@
It is possible to show that $\bar{\tau}_{t}^{x}=\tau_{t}^{x}$ a.s., where $\tau_{t}^{x}$ is the smallest optimal stopping time for the original problem defined in Corollary 3.1. A comparison of the three stopping times is given below.
@@
(1) By the definition of $\tau_{t}^{x}$ and the reward function $\varphi$, we have $v(\tau_{t}^{x}(\omega),F^{t,x}_{\tau_{t}^{x}(\omega)}(\omega))=g(\tau_{t}^{x}(\omega),F^{t,x}_{\tau_{t}^{x}(\omega)}(\omega))F^{t,x}_{\tau_{t}^{x}(\omega)}(\omega)$}, and if $\widetilde{\tau}_{t}^{x}(\omega)=\tau_{t}^{x}(\omega)$, it follows that $\widetilde{\tau}_{t}^{x}(\omega)=\tau_{t}^{x}(\omega)=\bar{\tau}_{t}^{x}(\omega).$} From the definition of $\tau_{t}^{x}(\omega)$, we know that $v(s,F^{t,x}_{s}(\omega))>g(s,F^{t,x}_{s}(\omega))F^{t,x}_{s}(\omega)$} for all $t\leq s<\tau^{x}_{t}(\omega).$} Thus, for $\widetilde{\tau}_{t}^{x}(\omega)$ to be equal to $\tau_{t}^{x}(\omega)$, we need $v(s,F^{t,x}_{s}(\omega))>h(s,F^{t,x}_{s}(\omega))$}, for all $t\leq s<\tau^{x}_{t}(\omega)$}. Now, suppose there exists some $s$ in the interval $t,\tau_{t}^{x}(\omega))$ such that $v(s,F^{t,x}_{s}(\omega))=h(s,F^{t,x}_{s}(\omega))$}. In this case, the two stopping times $\tau_{t}^{x}(\omega)$ and $\bar{\tau}_{t}^{x}(\omega)$ would differ. However, the equality $v(s,F^{t,x}_{s}(\omega))=h(s,F^{t,x}_{s}(\omega))$} would imply that the probability that the contract is exercised before maturity is nil, which contradicts our initial assumption that $\tau_{t}^{x}(\omega)<T$. Therefore, it must hold that $v(s,F^{t,x}_{s}(\omega))>h(s,F^{t,x}_{s}(\omega))$} for all $t\leq s<\tau^{x}_{t}(\omega)$}, which leads to $\widetilde{\tau}_{t}^{x}(\omega)=\tau_{t}^{x}(\omega)=\bar{\tau}_{t}^{x}(\omega).$
@@
The intuition behind the equality between $\tau^{x}_{t}$ and $\bar{\tau}^{x}_{t}$ is that $\bar{\tau}^{x}_{t}$ depends on why the contract is surrendered in the problem with the continuous reward. Indeed, the strategy underlying $\widetilde{\tau}^{x}_{t}$ is to stop if the contract value is equal to either the surrender benefit or the expected present value (EPV) of the maturity benefit. If the reason to stop is equality with the surrender benefit, then this optimal stopping time coincides with the strategy underlying the problem with the discontinuous reward function and all three stopping times are equal. However, if the reason to stop is that the value of the contract is equal to the EPV of the maturity benefit, then $\bar{\tau}^{x}_{t}$ is set to maturity $T$. In that case, it is also optimal to wait until maturity in the original problem, since the contract being worth the EPV of the maturity benefit means that $T$ is the optimal stopping time.
@@
In the next section, we present a condition on the fee and surrender charge functions under which $\widetilde{\tau}_{t}^{x}=\tau_{t}^{x}$ a.s. This result is particularly interesting since it provides a condition under which the two optimal stopping problems defined in (7) and (13) are equivalent. That is, the two problems have the same value function (Theorem 3.4), surrender region (Proposition 4.7), and optimal stopping strategy (Corollary 4.3).
@@ 4. Analytical study of the value function
The representation of the value function in terms of a continuous reward function simplifies the analysis of its regularity, which is presented in Section 4.1. However, additional work is required to assess its smoothness, which we discuss in Proposition 4.1 below. In Section 4.2, we establish the relationship between $v$, a free boundary value problem, and a variational inequality. This allows us to derive, in Section 4.3, two other representations for the value function: the surrender premium representation which is analogous to the exercise premium representation in the American option pricing terminology, and the continuation premium representation. This representation is new to the literature on VA and American option pricing and leads to the characterization of the (non-)emptiness of the surrender region in Section 4.4. Section 4.5 presents a condition under which the optimal stopping problem with discontinuous reward function defined in (7) is equivalent to the one with the continuous reward function, in (13).
Assumptions 4.1 and 4.2 below are needed for several of the results of this section to hold. In Assumption 4.1 below, $\mu:[0,T]\times\mathbb{R}_{+}\rightarrow\mathbb{R}$} denotes the drift term of the account value process (4), that is, $\mu(t,x)=(r-c(t,x))x$.
@@
    For all $x$, the function $t\mapsto c(t,x)$ is locally Hölder continuous.
@@ 4.1. Elementary Properties of the Value Function
The value function $v$ is continuous on $0,T]\times\mathbb{R}_{+}$.
@@
For every $(t,\,x)\in[0,\,T]\times\mathbb{R}_{+}$, the value function $v$ satisfies the following properties:
@@
The properties given in Lemma 4.1 follow from the definition of the contract. They do not require Assumptions 4.1 and 4.2 to be satisfied.
@@
Proposition 4.1 below concerns the smoothness of the value function. Our proof, which can be found in the Appendix, presents non-trivial differences compared to the American put case, which are due to the time-dependence of the value function.
@@ 4.2. Free Boundary Value Problem and Variational Inequality
The continuous reward representation of Theorem 3.4 and the previously established continuity of the value function allows us to apply classical results from optimal stopping theory. In this section, we use this strategy to establish the relationship between the value function $v$ associated with the original optimal stopping problem (7), a free boundary value problem, and a variational inequality. Thus, the proofs of the results presented in this section are kept short; the reader is referred to the literature for more details.
The value function is supported on $0,T]\times\mathbb{R}_{+}$, which is separated into the surrender region $\mathcal{S}$ and the continuation region $\mathcal{C}$, defined by
@@
respectively. It follows that $\mathcal{C}\cup\mathcal{S}=[0,T)\times\mathbb{R}_{+}$, since $v(t,x)\geq\varphi(t,x)$ for all $(t,x)\in[0,T]\times\mathbb{R}_{+}$. Furthermore, since $v$ and $\varphi$ are continuous on $0,T)\times\mathbb{R}_{+}$ (see Theorem 4.1), the continuation region is open and the surrender region is closed. The exercise (or surrender) boundary is the boundary $\partial\mathcal{C}$ of $\mathcal{C}$.
To express the value function as the solution of a free boundary problem, we define the second-order differential operator $\mathcal{L}_{t}$, for $0\leq t\leq T$, by
@@
Note that $L$ is the term on the left-hand side of (9). We give this general definition of $L$ since many of our results apply in a general setting, when the fee and surrender charge function depend on both $t$ and $x$. However, under Assumption 4.2, which is assumed to hold throughout Section 4, $L$ becomes
(19) $L(t)=g_{t}(t)-c(t)g(t).$
The next results confirm that $v(t,x)$ can be expressed as the solution to a free boundary problem.
@@
The value function $v$ solves the boundary value problem
@@
In particular, the function $v\in C^{1,2}$ on $\mathcal{C}$.
This result follows from the theory on optimal stopping, see for example Chapter 3 of Peskir and Shiryaev (2006).
Once continuity of the value function is verified (Theorem 4.1), the optimal stopping time is shown to exist (Corollary 3.1), and the martingale property of the Snell envelope of the discounted reward process is proven (see Theorem 3.3 (ii), the connection between boundary value and optimal stopping problems can be established assuming enough regularity of the coefficients in (4). A sufficient condition for the results of Friedman (1964) to hold is that the drift term in (4) is locally Hölder continuous, which is why Assumption 4.1 is needed. The proof makes use of standard partial differential equation (PDE) results in solving the Dirichlet (or the first initial-boundary value) problem, see for instance Friedman (1964), Theorem 3.4.9.
The main difficulty in establishing the free boundary value problem resides in showing that the value function is continuous (Theorem 4.1), which implies that the continuation region is open. Then, the Dirichlet problem (or the first initial-boundary value problem) can be posed in an open subset (typically a ball or a rectangle) of the continuation region. Since the boundary of the defined open subsets is sufficiently regular, standard PDE results guarantee that the Dirichlet problem admits a unique solution. Finally, probabilistic arguments are used to identify this solution as the value function.
For details of the proof of a similar result in the context of the American put in the Black-Scholes setting, see Jacka (1991), Proposition 2.6 or Karatzas and Shreve (1998), Theorem 2.7.7, among others. For more general results, see Peskir and Shiryaev (2006), Sections 3.7.1 and 4.8.2 or Jacka and Lynn (1992), Proposition 3.1.
@@
For the next remark, we recall that for an open set $O\subset\mathbb{R}^{d}$, $d\in\mathbb{N}$, and for $1\leq p<\infty$, $k\in\mathbb{N}$, the Sobolev space $\mathcal{W}^{k,p}(O)$ contains the $L^{p}(O)$ functions whose first $k$ weak partial derivatives have finite $L^{p}$-norm (see Evans (2010), Chapter 5 for more details). We further write that a function $f\in\mathcal{W}_{\text{loc}}^{k,p}(O)$ if $f\in\mathcal{W}^{k,p}(O_{1})$ for every bounded domain $O_{1}$ with $\bar{O}_{1}\subseteq O$.
Remark 4.8.
It is possible to show that $v$ is the unique solution to (20) among the functions $f\in\mathcal{W}_{loc}^{2,1}((0,T)\times\mathbb{R}_{+})$ satisfying $f(t,x)\leq G+x$ for all $(t,x)\in[0,T]\times\mathbb{R}_{+}$, see for example Section 8.2 and Chapter 11 of Pascucci (2011). Indeed, under Assumption 4.1, all the coefficients of $\mathcal{L}_{t}$ are Hölder continuous, so that Hypotheses 8.1 and 8.3 of Pascucci (2011) are satisfied and their Theorem 9.48 holds. In particular, this means that the weak derivatives $v_{t}$, $v_{x}$ and $v_{xx}$ are well-defined and integrable on $(0,T)\times\mathbb{R}_{+}$.
Assumption 4.1 does not give much information on the shape of the surrender region. However, we show in Section 4.4 (Theorem 4.3) that under further conditions on the fee and the surrender charge functions, the surrender region can be expressed in terms of the optimal surrender boundary $b:[0,T]\mapsto\mathbb{R}_{+}$. Then, the free boundary value problem can be stated more explicitly.
@@
If $L(t)<0$ for all $t\in[0,T)$, the value function $v$ solves the boundary value problem
@@
with $b(t):=\inf\{x\in\mathbb{R}_{+}:(t,x)\in\mathcal{S}\}$. In particular, the function $v$ is $C^{1,2}$ on $\mathcal{C}$.
@@
For each $t\in[0,T)$, $L(t)<0$ is equivalent to $\frac{\,\mathrm{d}\ln g(t)}{\,\mathrm{d}t}<c(t)$. From Gronwall’s inequality, this condition is satisfied if $g(t)<g(0)e^{\int_{0}^{T}c(s)\,\,\mathrm{d}s}$ for all $t\in[0,T]$.
@@
We know from Theorem 4.2 that $v\in C^{1,2}$ on $\mathcal{C}$, and from the definition of the value function on the surrender region that $v\in C^{1,2}$ on $\text{int}(\mathcal{S})$. It is known from the literature on optimal stopping that continuity of the second spatial derivative usually fails at the boundary $\partial\mathcal{C}$. We also have from Remark 4.8 that $v\in\mathcal{W}^{2,1}_{\text{loc}}((0,T)\times\mathbb{R}_{+})$. Then, by the Sobolev-Morrey embedding theorem (see for example Pascucci (2011), Theorem 8.16), $v$ and $v_{x}$ are Hölder-continuous of exponent $\alpha\in(0,1)$ on $0,T]\times\mathbb{R}_{+}$.
Next, we show that $v$ solves a variational inequality.
@@
The value function $v$ defined in (7) is a solution to the variational inequality
(22) $\max\left\{\mathcal{L}_{t}v+v_{t}-rv,\varphi-v\right\}=0,$
with terminal condition $v(T,x)=\max(G,x)$ and all derivatives in the weak sense.
Proof.
Since $v\in\mathcal{W}_{\text{loc}}^{2,1}((0,T)\times\mathbb{R}_{+})$ (Remark 4.8), the (weak) partial derivatives in (22) are well-defined. The Snell envelope $J=\{J_{t}\}_{0\leq t\leq T}$, with $J_{t}:=e^{-rt}v(t,\,F_{t})$ of the discounted reward process being a supermartingale (Theorem 3.3(i)),
(23) $\mathcal{L}_{t}v+v_{t}-rv\leq 0,$
with equality on $\mathcal{C}$ (Theorem 4.2). The result follows since $v(t,x)\geq\varphi(t,x)$ (Lemma 4.1) with equality on $\mathcal{S}$. ∎
@@
Proposition 4.2 holds as long as Assumption 4.1 is satisfied.
@@
The partial derivatives (in the sense of distribution) $v_{t}$, $v_{x}$ and $v_{xx}$ are locally bounded, and $v_{x}$ is continuous on $0,T)\times\mathbb{R}_{+}$.
@@
From Proposition 4.1, we have that the first order (weak) derivatives of the value function in $t$ and in $x$ are locally bounded. Using the convexity of $x\mapsto v(t,x)$ (Lemma 4.2), the local boundedness of the first order derivatives and (23), we can show that the second order derivative in $x$ is also locally bounded (see for instance Lamberton (1998), Theorem 10.3.8 or Jaillet et al. (1990), Theorem 3.6 for details). This last result is also known as the smooth fit condition.
The continuity of $v_{x}$ stems from Theorem 4.2 and the Sobolev-Morrey embedding theorem, as explained in Remark 4.10. ∎